# MagicClay: Sculpting Meshes With Generative Neural Fields

### v1

@@ 4.5. Ablations
We perform several ablations to justify the design of our system.
No face supersampling for mesh colors
Figure 6 illustrates the need for the mesh-driven super-sampling of the appearance network based on Mesh Colors schema (Sec 3.1). We optimize our representation with respect to a target image (first column), either sampling colors per-vertex (second column), or using our adaptive sampling using MeshColors (third column). Note that without MeshColors sampling a much higher resolution is needed, which leads to poorer reconstruction and an over-tesellated mesh.
Figure 6. Ablation: no color supersampling. Using mesh-guided supersampling in conjunction with the appearance network allows to decouple geometry and appearance. Using this approach (top right), large faces are used for the mesh, even though the rendering still presents high frequency colors (bottom). When using a color-per-vertex scheme (top middle), significantly larger mesh resolution is required to achieve similar appearance.
@@
We remove the mechanism for enforcing localization via fixing non-selected surface regions and nearby SDF values as discussed in Sec. 3.3. In Figure 7 we show that without this feature, the shapes undergo unintended global changes, potentially erasing the original shape.
Figure 7. Ablation: not enforcing localization. Without localization and freeze losses we observe that shape changes can propagate beyond the user-highlighted area, potentially destroying the original content. Here armadillo was erased by “angel wings.”
@@
The topological updates (Sec. 3.4) during the optimization significantly improve the results as they allow to gradually add resolution. Optimizing a fixed-resolution mesh would either result in a shape that only marginally differs from input if the initial resolution is too high (Fig 8, left), or lacks fine details if the initial resolution is too coarse (Fig 8, right).
Figure 8. Ablation: no topology updates. Optimizing the mesh without topology update results in the final generated object being limited by the initial resolution. Left When starting with a fine mesh the optimization will often get stuck since each vertex has tiny effect on the objective. Right When starting from a coarse mesh, no fine details can be created.
@@ 5. Conclusion
We presented MagicClay, a generative sculpting tool, backed by our new hybrid SDF and mesh representation. We demonstrated the importance of the hybrid representation through careful analysis and baselines. Key to the success of the generative process is our new rendering strategy that leverages the mesh part of the hybrid representation to localize ray sampled in the volumetric rendering of the SDF. We believe MagicClay is an important step towards turning the recent advancements in text-to-image-from-scratch into an actual modeling tool usable by artists in an iterative workflow.
@@
Our method is inherently constrained by the quality of the SDS gradients. This stems from the inability of the current generative model to generate consistent multi-view images during score-distillation sampling. Each view tracts the optimization in a different direction which results in a noisy process, rendering the emergence of fine details more difficult. That said, MagicClay employs current text-to-shape methods without adaptation, and can easily integrate and benefit from the unavoidable further development of the field. Second, MagicClay is not interactive, as running  MagicClay takes 1̃ hr per prompt on an A100 GPU. We know from the NeRF literature that a couple of dozens of multi-view images are sufficient to reconstruct 3D shapes very efficiently. Hence, we believe that as diffusion models become better at generating consistent images, the number of iterations required by  MagicClay will drop significantly, making it much faster.
@@
We see several venues for future research. First, we see opportunities to leverage inpainting and depth-conditioned diffusion models. Indeed, this generative process transforms the full object in each rendering, whereas it is clear that some part of the generated image should stay the same as the 3D edit is localized. We think that leveraging this insight would reduce the amount of noise in the 3D optimization helping the optimization reach a better geometry and faster. Second, we would like to explore using image targets with our system. We believe this could help making edits more specific, and also could allow more control, where users can highlight which part of the image should affect the generated shape. Finally, embellishing the surface representation to capture high resolution geometric details (such as normal and bump maps) and connecting it to SDF representation with appropriate consistency losses, can further improve the quality of the resulting shapes.

### v2

@@ 4.5. Ablations
512x512 SDF renderings
We show in Figure 6 and Table 1 that high-resolution SDF rendering significantly affects generated shape quality. This shows that the mesh part of the hybrid helping to accelerate SDF rendering plays a role in quality.
Initial editing.
In our experiments, only significant differences in the initial sculpt lead to different generated results. Typically, performing an initial sculpt does impact the result compared to just selecting a region with no sculpting. However, two initial sculpts that are similar will lead to the same final results (see Figure 10).
@@
We remove the mechanism for enforcing localization via fixing non-selected surface regions and nearby SDF values as discussed in Sec. 3.3. In Figure 11, The shapes undergo unintended global changes, potentially erasing the original shape.
@@
The topological updates (Sec. 3.4) allow to add resolution gradually. Optimizing a fixed-resolution mesh would either result in a shape that only marginally differs from the input if the initial resolution is too high or lacks fine details if the initial resolution is too coarse (Figure 5). These results complement, in a generative setting, the experiments of [Palfinger, 2022] [^1] We refer the reader to the videos in the Readme of their GitHub repository, performed with ground truth multi-view supervision which similarly shows that optimizing very high-poly meshes leads to local minima.
@@ 5. Conclusion
We presented MagicClay, a generative sculpting tool backed by our new hybrid SDF and mesh representation. Key to the success of the generative process is our new rendering strategy that leverages the mesh part of the hybrid representation to localize ray sampled in the volumetric rendering of the SDF. We believe MagicClay is an important step towards turning the recent advancements in text-to-image-from-scratch into an actual modeling tool usable by artists in an iterative workflow.
@@
Our method is inherently constrained by the quality of the SDS gradients. Each view tracts the optimization in a different direction adding noise and rendering the emergence of fine details more difficult. Second, MagicClay is not interactive, e.g. running  MagicClay takes 1 hour per prompt on an A100 GPU, with The bottleneck being the SDS loss
@@
We see opportunities to leverage inpainting and depth-conditioned diffusion models to speed up the convergence of SDS. Indeed, this generative process transforms the full object in each rendering, whereas it is clear that some parts of the generated image should stay the same as the 3D edit is localized. We think that leveraging this insight would reduce the amount of noise inherent in SDS.